You are performing 7 independent Bernoulli trials with p = 0.2 and q = 0.8. Calculate the probability of the stated outcome. Check your answer using technology. (Round your answer to five decimal places.) 


P(X≥3) = _______ 
Answer #1

n = 7 , p = 0.2

P(X) = ncx px ( 1 - p)n-x

So,

P( X >= 3) = 1 - P( X <= 2)

= 1 - [ P (x = 0) + P( X = 1) + P (X = 2) ]

= 1 - [ 7C0 0.20 0.87 + 7C1 0.21 0.86 + 7C2 0.22 0.86 ]

= 0.14803